# Bootstrapping Estimators based on the Block Maxima Method

### v1

@@ 9 Proofs
By Conditions 2.2(a) and 10.1 we have $r_{n3}=O(1/m^{\ast})=o(1)$. Furthermore, by Condition 10.1 and Lemma 3.11 from Dehling and Philipp, (2002) with $s=r=1/(2+\nu)$ and $t=\nu/(2+\nu)$, we have $\sup_{d\geq 2}|\operatorname{Cov}(S_{n,1}^{-},S_{n,1+d}^{-})|\lesssim 10k(m^{\ast})^{-1}\alpha^{\nu/(2+\nu)}(kr)$, whence $r_{n4}\lesssim r^{-1}n(m^{\ast})^{-2}\alpha^{\nu/(2+\nu)}(kr)=o((m^{*})^{-2})=o(1)$ by Condition 2.2 and the fact that $2/\nu<\omega$. Put together, in view of $\operatorname{E}[S_{n,j}^{-}]=0$, these convergences imply $p^{-1/2}\sum_{j=1}^{p}S_{n,j}^{-}=o_{\mathbb{P}}(1).$
As there is a lag of at least $kr$ between the observations making up $S_{n,j_{1}}^{+}$ and $S_{n,j_{2}}^{+}$ for $j_{1}\neq j_{2}$ a standard argument involving characteristic functions, a complex-valued version of Lemma 3.9 in Dehling and Philipp, (2002), Condition 2.2 (b) and Levy’s Continuity Theorem allows for assuming that the $(S_{n,j}^{+})_{j}$ are independent. Hence, we may subsequently apply Lyapunov’s central limit theorem. To verify its conditions, note that, since $\operatorname{Var}(p^{-1/2}\sum_{j=1}^{p}S_{n,j}^{-})=o(1)$ and by using (9.1) and (9.7), $\operatorname{Var}(p^{-1/2}\sum_{j=1}^{p}S_{n,j}^{+})=\operatorname{Var}(\bar{\mathbb{G}}_{n,r}h)+o(1)=\Sigma_{h}^{({\operatorname{sb}})}+o(1).$ Hence,
@@
by Condition 10.1 and the choice of $m^{\ast}.$ An application of Lyapunov’s central limit theorem implies the assertion.
@@
$\tilde{\mathbb{G}}_{n,r}^{({\operatorname{cb}})}h=\bar{\mathbb{G}}_{n,r}^{({\operatorname{cb}})}h+\sqrt{\frac{n}{r}}(P_{n,r}^{({\operatorname{cb}})}-P_{r})h,\quad\mathbb{G}_{n,r}^{({\operatorname{cb}})}h=\bar{\mathbb{G}}_{n,r}^{({\operatorname{cb}})}h+\sqrt{\frac{n}{r}}(P_{n,r}^{({\operatorname{cb}})}-P_{r}+P_{r}-P)h.$
@@
$(P_{n,r}^{({\operatorname{cb}})}-P_{r})h=$ $\frac{1}{m(k)}\sum_{i=1}^{m(k)}\frac{1}{kr}\sum_{s=1}^{kr}\operatorname{E}[h(Z_{r,(i-1)kr+s})]-P_{r}h$
@@
Condition 10.3 then gives $\sqrt{n/r}(P_{n,r}^{({\operatorname{cb}})}-P_{r})h\to k^{-1}(D_{h,k}+E_{h}).$ Furthermore, Condition 10.2 immediately implies $\sqrt{n/r}(P_{r}-P)h\to B_{h}.$ Combined with the first assertion and the penultimate display, we obtain the additional two claimed weak convergences. ∎
@@
The second and third assertion follow immediately from the first one and the triangular inequality. For the proof of the first one, we omit the upper index ${\operatorname{cb}}$. By Theorem 2.3 (for ${\operatorname{mb}}={\operatorname{sb}}$) or Theorem 3.2 (for ${\operatorname{mb}}={\operatorname{cb}}$) and Lemma 2.3 from Bücher and Kojadinovic, (2019), it is sufficient to show that, unconditionally,
$(\hat{\mathbb{G}}_{n,r}^{[1]}\bm{h},\hat{\mathbb{G}}_{n,r}^{[2]}\bm{h})\rightsquigarrow\mathcal{N}_{q}(\bm{0},\Sigma^{({\operatorname{sb}})}_{h})\otimes\mathcal{N}_{q}(\bm{0},\Sigma^{({\operatorname{sb}})}_{h}),$
@@
$T_{n}$ $:=\sum_{b=1}^{2}a_{b}\sqrt{\frac{n}{r}}\frac{1}{n}\sum_{i=1}^{m(k)}\big{(}W_{m(k),i}^{[b]}-1\big{)}\sum_{s\in I_{kr,i}}h(\bm{Z}_{r,s})\rightsquigarrow\mathcal{N}_{2}\big{(}0,(a_{1}^{2}+a_{2}^{2})\Sigma_{h}^{({\operatorname{sb}})}\big{)}$ (9.8)
@@
$=\sum_{b=1}^{2}a_{b}\sqrt{\frac{n}{r}}\frac{1}{m(k)}\sum_{i=1}^{m(k)}\big{(}W_{m(k),i}^{[b]}-1\big{)}D_{n,i}$ (9.9)
with $D_{n,i}$ from (9.2). The subsequent proof strategy, known as Poissonization, consists of removing the dependence of the multinomial multipliers by introducing rowwise i.i.d. multiplier sequences that approximate the multinomial multipliers. More precisely, we employ the construction from Lemma 11.3: for each fixed $n\in\mathbb{N}$, let $(\bm{U}_{j,m(k)}^{[b]})_{j\in\mathbb{N},b\in\{1,2\}}$ be i.i.d. multinomial vectors with 1 trial and $m(k)$ classes, with class probabilities $1/m(k)$ for each class, and independent of $(\bm{X}_{t})_{t\in\mathbb{Z}}$. We may then assume that
@@
By Lemma 11.3 the random vectors $\tilde{\bm{W}}_{m(k)}^{[1]},\tilde{\bm{W}}_{m(k)}^{[2]}$ are i.i.d. $\mathrm{Poisson}(1)^{\otimes m(k)}$ distributed. Then, in view of (9), we may write $T_{n}=\tilde{T}_{n}+a_{1}R_{n1}+a_{2}R_{n2}$, where
@@
We may apply Lemma 11.2 to obtain that $\tilde{T}_{n}\rightsquigarrow\mathcal{N}(0,(a_{1}^{2}+a_{2}^{2})\Sigma_{h}^{({\operatorname{sb}})})$. For the proof of (9.8), it hence remains to verify that $R_{nb}=o_{\mathbb{P}}(1)$ for $b\in\{1,2\}$. We will suppress the upper index $[b]$ in the following. First write
@@
Therefore, since $|N_{m(k)}-m(k)|/\sqrt{m(k)}=O_{\mathbb{P}}(1)$ by the Central Limit Theorem, the proof of (9.8) is finished if we show that $\tilde{R}_{nb}=o_{\mathbb{P}}(1)$. For that purpose note that, for any fixed $\varepsilon>0$,
@@
where we used $\sum_{i=1}^{m(k)}H_{n,i}\leq 1$. Since $\max_{j=1}^{m(k)}H_{n,j}=o_{\mathbb{P}}(1)$ and by the (proof of the) three assertions in (9.4), we obtain that the expression on the right-hand side of the previous display is $o_{\mathbb{P}}(1)$. Hence, writing $\mathbb{P}(|\tilde{R}_{nb}|>\varepsilon)=\operatorname{E}\big{[}\mathbb{P}\big{(}|\tilde{R}_{nb}|>\varepsilon\mid(H_{n,i})_{i}\big{)}\big{]}$ and invoking the Dominated Convergence Theorem for convergence in probability, we obtain $\tilde{R}_{nb}=o_{\mathbb{P}}(1)$ and the proof is finished. ∎
Proof for Remark 4.3.
We sketch the proof of the conditional weak convergence in (4.2) provided that Conditions 2.1, 2.2 and  10.1(a) are met. Repeating the arguments from the proof of Theorem 4.1, the assertion follows from unconditional weak convergence of
$\tilde{\mathbb{G}}_{n,r}^{({\operatorname{sb}}),*}h=\sqrt{\frac{n}{r}}\frac{1}{m(k)}\sum_{i=1}^{m(k)}Y_{n,i}D_{n,i},$
where $(Y_{n,i})_{i}$ are iid with expectation 0 and variance 1, and where $D_{n,i}$ are defined as in (9.2), but with ${\operatorname{sb}}$ instead of ${\operatorname{cb}}$. We only provide the proof for the convergence of the variance of $\tilde{\mathbb{G}}_{n,r}^{({\operatorname{sb}}),*}h$ to $\Sigma_{h}^{(k)}$; the remaining arguments are then the same as in the proof of Lemma 11.2.
Since $(Y_{n,i}D_{n,i})_{i}$ in uncorrelated, we obtain
$\operatorname{Var}(\tilde{\mathbb{G}}_{n,r}^{({\operatorname{sb}}),*)}h)=k\operatorname{Var}(D_{n,1})=\frac{1}{kr^{2}}\sum_{s=1}^{kr}\sum_{t=1}^{kr}\operatorname{Cov}(h(\bm{Z}_{r,s}^{({\operatorname{sb}})}),h(\bm{Z}_{r,t}^{({\operatorname{sb}})}))$
Using the notation after (9.4), we obtain
$\lim_{n\to\infty}\operatorname{Var}(\tilde{\mathbb{G}}_{n,r}^{({\operatorname{sb}}),*)}h)=\frac{1}{k}\int_{0}^{k}\int_{0}^{k}g(\xi,\xi^{\prime})\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi.$
We start by considering the case $k\geq 2$. Then, by the calculation in (9) but with $g_{(k)}$ replaced by $g$,
where
$I_{1}$ $=(k-1)\Sigma_{h}^{({\operatorname{sb}})}/2+\int_{0}^{1}\int_{0}^{y}u(x)\,\mathrm{d}x\,\mathrm{d}y,\qquad I_{2}=\int_{0}^{1}\int_{k-1+\xi}^{k}u(\xi^{\prime}-\xi)\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi=0.$
The first integral over $u$ can be rewritten as
$\int_{0}^{1}\int_{0}^{y}u(x)\,\mathrm{d}x\,\mathrm{d}y=\int_{0}^{1}\int_{x}^{1}u(x)\,\mathrm{d}y\,\mathrm{d}x=\int_{0}^{1}(1-x)u(x)\,\mathrm{d}x=\Sigma_{h}^{({\operatorname{sb}})}/2-\int_{0}^{1}xu(x)\,\mathrm{d}x,$ (9.10)
so that
$I_{1}=k\Sigma_{h}^{({\operatorname{sb}})}/2-\int_{0}^{1}xu(x)\,\mathrm{d}x.$
$\operatorname{Var}(\tilde{\mathbb{G}}_{n,r}^{({\operatorname{sb}}),*}h)=k\operatorname{Var}(D_{n,1})=\Sigma_{h}^{({\operatorname{sb}})}-\frac{2}{k}\int_{0}^{1}xu(x)\,\mathrm{d}x+o(1)$
as asserted.
which implies the assertion by (9.10). ∎
Proof of Theorem 5.5.
@@
The results regarding $(\hat{\alpha}_{n}^{({\operatorname{cb}})},\hat{\sigma}_{n}^{({\operatorname{cb}})})^{\top}$ and $(\hat{\alpha}_{n}^{({\operatorname{cb}}),*},\hat{\sigma}_{n}^{({\operatorname{cb}}),*})^{\top}$ follow from an application of Theorem 11.6 and Theorem 11.7, respectively.
@@
For the former we need to show that, with $v_{n}=\sqrt{n/r},\omega_{n}=n$ and $X_{n,s}=M_{r,s}^{({\operatorname{cb}})}\vee c$, the “no-ties-condition” in Equation (11.4) is met, and that the three convergences in Condition 11.5 are satisfied. Equation (11.4) follows from $\mathbb{P}(X_{n,1}=\dots=X_{n,n})\leq\mathbb{P}(X_{n,1}=X_{n,r+1})=\mathbb{P}((M_{r,1}\vee c)/\sigma_{r}=(M_{r,r+1}\vee c)/\sigma_{r})$, which converges to zero by the Portmanteau theorem, observing that $(M_{r,1}\vee c)/\sigma_{r},(M_{r,r+1}\vee c)/\sigma_{r})$ weakly converges to the product of two independent Fréchet$(\alpha_{0},1)$ random variables by Lemma 5.1 in Bücher and Segers, 2018a . Condition 11.5(iii) is a consequence of Condition 5.4. Condition 11.5(ii) follows from a straightforward modification of Theorem 3.2 to the Fréchet domain of attraction. Finally, for $\alpha_{+}$ chosen sufficiently close to $\alpha_{0}$, Condition 11.5(ii) follows from such a modification as well, following the argumentation on page 1116 in Bücher and Segers, 2018b .
@@
Next, consider the assertions regarding $(\hat{\alpha}_{n}^{({\operatorname{cb}}),*},\hat{\sigma}_{n}^{({\operatorname{cb}}),*})^{\top}$, which follow from Theorem 11.7 if we additionally show that the “no-ties-condition” in Equation (11.5) is met, that ^$\hat{\mathbb{P}}_{n,r}^{({\operatorname{cb}}),*}f=P_{\alpha_{0},1}f+o_{\mathbb{P}}(1)$ for all $f\in\mathcal{F}(\alpha_{-},\alpha_{+})$ and that (11.7) is met. The latter two assertions follow from similar arguments as given in the proof of Theorem 4.1 (adaptations to the Fréchet domain of attraction are needed), and it remains to show (11.5). For that purpose, let $I_{1},\dots,I_{m(k)}$ be iid uniformly distributed on $\{1,\dots,m(k)\}$ independent of $(X_{n,1},\ldots,X_{n,n})$, such that
@@
This probability is bounded by $\mathbb{P}(X_{n,1}=X_{n,r+1})$, which converges to zero as shown above when proving Equation (11.4). ∎
Proof of Corollary 5.6.
@@
Throughout, we omit the upper indexes $({\operatorname{sb}})$ and $({\operatorname{cb}})$ at $\hat{\sigma}_{n}$ and $\hat{\sigma}_{n}^{*}$, respectively. Define $S_{n}=\sqrt{\tfrac{n}{r}}(\hat{\sigma}_{n}/\sigma_{r_{n}}-1)$ and $S_{n}^{*}=\sqrt{\tfrac{n}{r}}(\hat{\sigma}_{n}^{*}-\hat{\sigma}_{n})/\sigma_{r_{n}}$, and note that, with $M_{2}(\alpha_{0})$ denoting the second row of $M(\alpha_{0})$,
$S_{n}=M_{2}(\alpha_{0})\mathbb{G}_{n,r}(f_{1},f_{2},f_{3})+o_{\mathbb{P}}(1),\qquad S_{n}^{*}=\frac{\hat{\sigma}_{n}}{\sigma_{r_{n}}}\sqrt{\frac{n}{r}}(\hat{\sigma}_{n}^{*}/\hat{\sigma}_{n}-1)=M_{2}(\alpha_{0})\hat{\mathbb{G}}_{n,r}^{*}(f_{1},f_{2},f_{3})+o_{\mathbb{P}}(1)$
@@
by Theorem 5.5. Moreover, arguing as in the proof of that theorem, the weak limit of $S_{n}$ coincides with the conditional weak limit of $\hat{S}_{n}^{*}$ given the observations $\mathcal{X}_{n}$. Therefore, observing that $F_{S_{n}^{*}}(x)=\mathbb{P}(S_{n}^{*}\leq x\mid\mathcal{X}_{n})$ satisfies $F_{S_{n}^{*}}(x)=F_{\hat{\sigma}_{n}^{*}}(\hat{\sigma}_{n}+\sqrt{\tfrac{r}{n}}\cdot\sigma_{r_{n}}x)$ and hence

### v2

@@ Appendix B Proofs
By Conditions 2.2(a) and 9.1 we have $r_{n3}=O(1/m^{\ast})=o(1)$. Furthermore, by Condition 9.1 and Lemma 3.11 from DehPhi02 with $s=r=1/(2+\nu)$ and $t=\nu/(2+\nu)$, we have $\sup_{d\geq 2}|\operatorname{Cov}(S_{n,1}^{-},S_{n,1+d}^{-})|\lesssim 10k(m^{\ast})^{-1}\alpha^{\nu/(2+\nu)}(kr)$, whence $r_{n4}\lesssim r^{-1}n(m^{\ast})^{-2}\alpha^{\nu/(2+\nu)}(kr)=o((m^{*})^{-2})=o(1)$ by Condition 2.2 and the fact that $2/\nu<\omega$. Put together, in view of $\operatorname{E}[S_{n,j}^{-}]=0$, these convergences imply $p^{-1/2}\sum_{j=1}^{p}S_{n,j}^{-}=o_{\mathbb{P}}(1).$
As there is a lag of at least $kr$ between the observations making up $S_{n,j_{1}}^{+}$ and $S_{n,j_{2}}^{+}$ for $j_{1}\neq j_{2}$ a standard argument involving characteristic functions, a complex-valued version of Lemma 3.9 in DehPhi02, Condition 2.2 (b) and Levy’s Continuity Theorem allows for assuming that the $(S_{n,j}^{+})_{j}$ are independent. Hence, we may subsequently apply Lyapunov’s central limit theorem. To verify its conditions, note that, since $\operatorname{Var}(p^{-1/2}\sum_{j=1}^{p}S_{n,j}^{-})=o(1)$ and by using (B.1) and (B.7), $\operatorname{Var}(p^{-1/2}\sum_{j=1}^{p}S_{n,j}^{+})=\operatorname{Var}(\bar{\mathbb{G}}_{n,r}h)+o(1)=\Sigma_{h}^{[{\operatorname{s}}]}+o(1).$ Hence,
@@
by Condition 9.1 and the choice of $m^{\ast}.$ An application of Lyapunov’s central limit theorem implies the assertion.
@@
$\tilde{\mathbb{G}}_{n,r}^{[{\operatorname{c}}]}h=\bar{\mathbb{G}}_{n,r}^{[{\operatorname{c}}]}h+\sqrt{\frac{n}{r}}(P_{n,r}^{[{\operatorname{c}}]}-P_{r})h,\quad\mathbb{G}_{n,r}^{[{\operatorname{c}}]}h=\bar{\mathbb{G}}_{n,r}^{[{\operatorname{c}}]}h+\sqrt{\frac{n}{r}}(P_{n,r}^{[{\operatorname{c}}]}-P_{r}+P_{r}-P)h.$
@@
$(P_{n,r}^{[{\operatorname{c}}]}-P_{r})h=$ $\frac{1}{m(k)}\sum_{i=1}^{m(k)}\frac{1}{kr}\sum_{s=1}^{kr}\operatorname{E}[h(Z_{r,(i-1)kr+s})]-P_{r}h$
@@
Condition 9.3 then gives $\sqrt{n/r}(P_{n,r}^{[{\operatorname{c}}]}-P_{r})h\to k^{-1}(D_{h,k}+E_{h}).$ Furthermore, Condition 9.2 immediately implies $\sqrt{n/r}(P_{r}-P)h\to B_{h}.$ Combined with the first assertion and the penultimate display, we obtain the additional two claimed weak convergences. ∎
@@
The second and third assertion follow immediately from the first one and the triangular inequality. For the proof of the first one, we omit the upper index ${\operatorname{c}}$. By Theorem 2.3 (for ${\operatorname{m}}={\operatorname{s}}$) or Theorem 3.2 (for ${\operatorname{m}}={\operatorname{c}}$) and Lemma 2.3 from BucKoj2019note, it is sufficient to show that, unconditionally,
$(\hat{\mathbb{G}}_{n,r}^{[1]}\bm{h},\hat{\mathbb{G}}_{n,r}^{[2]}\bm{h})\rightsquigarrow\mathcal{N}_{q}(\bm{0},\Sigma^{[{\operatorname{s}}]}_{h})\otimes\mathcal{N}_{q}(\bm{0},\Sigma^{[{\operatorname{s}}]}_{h}),$
@@
$T_{n}$ $:=\sum_{b=1}^{2}a_{b}\sqrt{\frac{n}{r}}\frac{1}{n}\sum_{i=1}^{m(k)}\big{(}W_{m(k),i}^{[b]}-1\big{)}\sum_{s\in I_{kr,i}}h(\bm{Z}_{r,s})\rightsquigarrow\mathcal{N}_{2}\big{(}0,(a_{1}^{2}+a_{2}^{2})\Sigma_{h}^{[{\operatorname{s}}]}\big{)}$ (B.8)
@@
$=\sum_{b=1}^{2}a_{b}\sqrt{\frac{n}{r}}\frac{1}{m(k)}\sum_{i=1}^{m(k)}\big{(}W_{m(k),i}^{[b]}-1\big{)}D_{n,i}$ (B.9)
with $D_{n,i}$ from (B.2). The subsequent proof strategy, known as Poissonization, consists of removing the dependence of the multinomial multipliers by introducing rowwise i.i.d. multiplier sequences that approximate the multinomial multipliers. More precisely, we employ the construction from Lemma C.3: for each fixed $n\in\mathbb{N}$, let $(\bm{U}_{j,m(k)}^{[b]})_{j\in\mathbb{N},b\in\{1,2\}}$ be i.i.d. multinomial vectors with 1 trial and $m(k)$ classes, with class probabilities $1/m(k)$ for each class, and independent of $(\bm{X}_{t})_{t\in\mathbb{Z}}$. We may then assume that
@@
By Lemma C.3 the random vectors $\tilde{\bm{W}}_{m(k)}^{[1]},\tilde{\bm{W}}_{m(k)}^{[2]}$ are i.i.d. $\mathrm{Poisson}(1)^{\otimes m(k)}$ distributed. Then, in view of (B), we may write $T_{n}=\tilde{T}_{n}+a_{1}R_{n1}+a_{2}R_{n2}$, where
@@
We may apply Lemma C.2 to obtain that $\tilde{T}_{n}\rightsquigarrow\mathcal{N}(0,(a_{1}^{2}+a_{2}^{2})\Sigma_{h}^{[{\operatorname{s}}]})$. For the proof of (B.8), it hence remains to verify that $R_{nb}=o_{\mathbb{P}}(1)$ for $b\in\{1,2\}$. We will suppress the upper index $[b]$ in the following. First write
@@
Therefore, since $|N_{m(k)}-m(k)|/\sqrt{m(k)}=O_{\mathbb{P}}(1)$ by the Central Limit Theorem, the proof of (B.8) is finished if we show that $\tilde{R}_{nb}=o_{\mathbb{P}}(1)$. For that purpose note that, for any fixed $\varepsilon>0$,
@@
where we used $\sum_{i=1}^{m(k)}H_{n,i}\leq 1$. Since $\max_{j=1}^{m(k)}H_{n,j}=o_{\mathbb{P}}(1)$ and by the (proof of the) three assertions in (B.4), we obtain that the expression on the right-hand side of the previous display is $o_{\mathbb{P}}(1)$. Hence, writing $\mathbb{P}(|\tilde{R}_{nb}|>\varepsilon)=\operatorname{E}\big{[}\mathbb{P}\big{(}|\tilde{R}_{nb}|>\varepsilon\mid(H_{n,i})_{i}\big{)}\big{]}$ and invoking the Dominated Convergence Theorem for convergence in probability, we obtain $\tilde{R}_{nb}=o_{\mathbb{P}}(1)$ and the proof is finished. ∎
Proof of Proposition 4.4.
By Equations (2.5), (2.6), (2.8) and (4.3), (4.4) we have
$\hat{\bm{\theta}}_{n}^{[{\operatorname{m}}]}-{\bm{\theta}}_{r}$ $=A(\bm{a}_{r},\bm{b}_{r})\sqrt{\frac{r}{n}}\big{(}\bar{\mathbb{G}}_{n,r}^{[{\operatorname{m}}]}\bm{h}+o_{\mathbb{P}}(1)\big{)}=:\sqrt{\frac{r}{n}}A(\bm{a}_{r},\bm{b}_{r})S_{n},$
$\hat{\bm{\theta}}_{n}^{[{\operatorname{c}}],*}-\hat{\bm{\theta}}_{n}^{[{\operatorname{c}}]}$ $=A(\bm{a}_{r},\bm{b}_{r})\sqrt{\frac{r}{n}}\big{(}\hat{\mathbb{G}}_{n,r}^{({\operatorname{c}},\ast)}\bm{h}+o_{\mathbb{P}}(1)\big{)}=:\sqrt{\frac{r}{n}}A(\bm{a}_{r},\bm{b}_{r})S_{n}^{*}.$
By Theorems 2.3 and 3.2 we have $S_{n}\rightsquigarrow\mathcal{N}_{q}(\bm{0},\Sigma_{\bm{h}}^{[{\operatorname{s}}]})$. Theorem 4.1 implies $d_{K}(\mathcal{L}(S_{n}^{*}\mid\mathcal{X}_{n}),\mathcal{L}(S_{n}))=o_{\mathbb{P}}(1).$ The assertion then follows from Lemma C.9 of the supplement. ∎
Proof of Theorem A.5.
@@
The results regarding $(\hat{\alpha}_{n}^{[{\operatorname{c}}]},\hat{\sigma}_{n}^{[{\operatorname{c}}]})^{\top}$ and $(\hat{\alpha}_{n}^{[{\operatorname{c}}],*},\hat{\sigma}_{n}^{[{\operatorname{c}}],*})^{\top}$ follow from an application of Theorem C.6 and Theorem C.7, respectively.
@@
For the former we need to show that, with $v_{n}=\sqrt{n/r},\omega_{n}=n$ and $X_{n,s}=M_{r,s}^{[{\operatorname{c}}]}\vee c$, the “no-ties-condition” in Equation (C.4) is met, and that the three convergences in Condition C.5 are satisfied. Equation (C.4) follows from $\mathbb{P}(X_{n,1}=\dots=X_{n,n})\leq\mathbb{P}(X_{n,1}=X_{n,r+1})=\mathbb{P}((M_{r,1}\vee c)/\sigma_{r}=(M_{r,r+1}\vee c)/\sigma_{r})$, which converges to zero by the Portmanteau theorem, observing that $(M_{r,1}\vee c)/\sigma_{r},(M_{r,r+1}\vee c)/\sigma_{r})$ weakly converges to the product of two independent Fréchet$(\alpha_{0},1)$ random variables by Lemma 5.1 in BucSeg18-sl. Condition C.5(iii) is a consequence of Condition A.4. Condition C.5(ii) follows from a straightforward modification of Theorem 3.2 to the Fréchet domain of attraction. Finally, for $\alpha_{+}$ chosen sufficiently close to $\alpha_{0}$, Condition C.5(i) follows from such a modification as well, following the argumentation on page 1116 in BucSeg18a.
@@
Next, consider the assertions regarding $(\hat{\alpha}_{n}^{[{\operatorname{c}}],*},\hat{\sigma}_{n}^{[{\operatorname{c}}],*})^{\top}$, which follow from Theorem C.7 if we additionally show that the “no-ties-condition” in Equation (C.5) is met, that ^$\hat{\mathbb{P}}_{n,r}^{[{\operatorname{c}}],*}f=P_{\alpha_{0},1}f+o_{\mathbb{P}}(1)$ for all $f\in\mathcal{F}(\alpha_{-},\alpha_{+})$ and that (C.7) is met. The latter two assertions follow from similar arguments as given in the proof of Theorem 4.1 (adaptations to the Fréchet domain of attraction are needed), and it remains to show (C.5). For that purpose, let $I_{1},\dots,I_{m(k)}$ be iid uniformly distributed on $\{1,\dots,m(k)\}$ independent of $(X_{n,1},\ldots,X_{n,n})$, such that
@@
This probability is bounded by $\mathbb{P}(X_{n,1}=X_{n,r+1})$, which converges to zero as shown above when proving Equation (C.4). ∎
Proof of Corollary A.6.
@@
Define $S_{n}^{[{\operatorname{s}}]}=\sqrt{\tfrac{n}{r}}(\hat{\sigma}_{n}^{[{\operatorname{s}}]}/\sigma_{r_{n}}-1)$ and $S_{n}^{[{\operatorname{c}}],*}=\sqrt{\tfrac{n}{r}}(\hat{\sigma}_{n}^{[{\operatorname{c}}],*}-\hat{\sigma}_{n}^{[{\operatorname{c}}]})/\sigma_{r_{n}}$, and note that, with $M_{2}(\alpha_{0})$ denoting the second row of $M(\alpha_{0})$,
$S_{n}^{[{\operatorname{s}}]}=M_{2}(\alpha_{0})\mathbb{G}_{n,r}^{[{\operatorname{s}}]}(f_{1},f_{2},f_{3})+o_{\mathbb{P}}(1),\qquad S_{n}^{[{\operatorname{c}}],*}=\frac{\hat{\sigma}_{n}^{[{\operatorname{c}}]}}{\sigma_{r_{n}}}\sqrt{\frac{n}{r}}\Big{(}\frac{\hat{\sigma}_{n}^{[{\operatorname{c}}],*}}{\hat{\sigma}_{n}^{[{\operatorname{c}}]}}-1\Big{)}=M_{2}(\alpha_{0})\hat{\mathbb{G}}_{n,r}^{[{\operatorname{c}}],*}(f_{1},f_{2},f_{3})+o_{\mathbb{P}}(1)$
@@
by Theorem A.5. Moreover, arguing as in the proof of that theorem, the weak limit of $S_{n}^{[{\operatorname{s}}]}$ coincides with the conditional weak limit of $\hat{S}_{n}^{[{\operatorname{c}}],*}$ given the observations $\mathcal{X}_{n}$. Therefore, observing that $F_{S_{n}^{[{\operatorname{c}}],*}}(x)=\mathbb{P}(S_{n}^{[{\operatorname{c}}],*}\leq x\mid\mathcal{X}_{n})$ satisfies $F_{S_{n}^{[{\operatorname{c}}],*}}(x)=F_{\hat{\sigma}_{n}^{[{\operatorname{c}}],*}}(\hat{\sigma}_{n}^{[{\operatorname{c}}]}+\sqrt{\tfrac{r}{n}}\cdot\sigma_{r_{n}}x)$ and hence